OpenStudy (anonymous):

Solve for x and check. Enter only a number. 3^2x = 27

OpenStudy (anonymous):

rewrite 27 as a base 3 exponent 3^(2x)=3^3 take the log base 3 of each side log_3(3^2x)=log_3(3^3) 2xlog_3(3)=3log_3(3) log_3(3)=1 2x(1)=3(1) 2x=3 x=3/2

OpenStudy (anonymous):

Yes I agree with D above Also in this special case, one could also use the inspection as follows: \[3^{2x} = 27 \] \[=> 3^{2x} = 3^{3}\]=> 2x=3 =>x =3/2